A solid wood door 1.00 m wide and 2.00 m high is hinged along one side and has a total mass of 42.0 kg . Initially open and at rest, the door is struck at its center by a handful of sticky mud with mass 0.300 kg , traveling perpendicular to the door at 11.0 m/s just before impact. A: Find the final angular speed of the door.
Answer #1

Moment of inertia If = ML²/3 + m(L/2)² = 42kg * (1m)² / 3 + 0.300kg * (0.5m)²
If = 14kg·m² + 0.075kg·m²
If = 14.075 kg·m²

Initial angular momentum Li = mvr = 0.300kg * 11.0m/s * 0.5m = 1.65 kg·m²/s
Final angular momentum Lf = I*ω = 14.075 kg·m² * ω

Using Conservation of Momentum,
1.65 =  14.075 * ω
ω = 0.117 rad/s
final angular speed of the door, ω = 0.117 rad/s